OpenStudy (anonymous):

What is the total surface area of the tent in square feet? Assume no material is used on the bottom of the tent.

OpenStudy (cwrw238):

you have 4 walss 2*(7*3)and 2*(6*3) two parts of the roof = 2*5*7 and 2 triangular parts = 2*3*4 sq ft calculate these and add them up to get total area

OpenStudy (anonymous):

172 ?

OpenStudy (cwrw238):

hmm 42 + 36 + 70 + 24 = 172 dead right!

OpenStudy (anonymous):

MUCHO gracias !

OpenStudy (cwrw238):

nada - its that correct spanish?

OpenStudy (anonymous):

lol it should be im not spanish

OpenStudy (cwrw238):

oh ok - lol

OpenStudy (anonymous):

u think u can help w/ another problem ?